### v1

@@ 4.3 Multimodal Datasets
ChatGPT 69.417 - - 56.553 53.187 48.173 89.605 44.4410 84.975 40.771 65.484 59.9710 72.703 60.848 69.027 38.282 - 853.38
LLaMA1-7B 36.195 - - 67.231 60.721 38.261 82.012 34.262 75.621 15.501 58.533 46.433 65.677 50.404 58.994 2.663 - 692.47
LLaMA1-13B 43.336 - - 66.992 68.828 44.686 72.105 34.112 79.552 28.742 52.073 47.247 65.492 49.205 57.534 19.225 - 729.07
LLaMA2-7B 43.086 - - 66.991 69.222 40.284 67.681 26.381 77.301 16.781 58.533 46.602 67.192 52.001 56.331 3.597 - 691.95
LLaMA2-13B 55.787 - - 66.021 68.697 45.786 81.862 31.496 81.662 24.336 60.375 48.545 69.102 55.002 60.231 20.005 - 768.85
Mixtral-AWQ 66.633 - - 54.379 55.5910 43.941 87.925 41.237 79.305 38.277 55.451 60.212 70.561 58.005 64.387 37.039 - 812.88
Gemma 52.955 - - 67.728 61.619 43.102 81.657 36.117 77.052 25.067 54.295 52.435 70.945 57.801 60.072 30.312 - 771.09
Flan-T5-XXL 67.439 - - 64.814 66.014 49.563 89.6010 42.866 86.526 46.296 72.135 63.703 74.398 57.407 71.409 31.419 - 883.51
GPT-4V 79.8010 31.113 57.789 76.195 71.054 50.583 90.913 61.193 87.108 49.441 53.858 60.162 78.953 78.008 76.769 54.298 1057.16 968.27
Claude3-V 76.355 36.1110 53.337 80.954 69.085 46.1510 78.791 32.841 79.937 30.731 38.465 54.475 76.321 66.005 71.372 57.1410 948.02 858.58
Gemini-V 80.355 26.727 47.117 72.733 70.183 51.653 88.348 48.688 87.147 38.867 54.518 59.328 81.023 68.324 56.8310 41.414 973.17 899.34
OpenFlamingo 39.275 23.675 24.447 55.587 61.157 29.479 79.977 24.852 77.307 12.122 57.285 46.195 66.917 49.402 52.681 3.126 703.40 655.29
Fromage 34.517 21.947 20.337 29.856 28.192 22.761 57.197 19.152 47.412 11.042 19.966 27.316 35.106 37.602 40.687 0.161 453.18 410.91
LLaVA-v0-7B 41.105 19.945 24.854 69.423 65.426 30.448 74.692 30.039 74.657 18.129 35.103 44.573 73.483 39.805 43.2110 4.068 688.88 644.09
LLaVA-v0-13B 47.747 12.612 15.187 73.067 69.612 38.514 80.187 30.906 76.587 28.373 37.995 48.465 77.699 40.405 44.295 3.915 725.48 697.69
LLaVA-v1.6-7B 65.895 28.114 34.767 59.952 67.232 45.121 85.639 30.416 81.621 23.977 59.319 52.841 75.189 57.802 59.615 17.191 844.62 781.75
LLaVA-v1.6-13B 69.665 26.899 39.533 64.563 60.433 53.223 86.393 35.969 78.263 40.897 58.737 56.0810 77.9510 61.204 62.313 22.813 894.87 828.45
MiniGPT4 58.705 27.335 30.565 71.125 70.782 50.294 83.994 35.422 83.382 38.465 47.167 49.439 76.012 54.2010 57.493 2.816 837.13 779.24
mPLUG-Owl 37.937 26.173 26.226 51.946 50.366 33.371 68.751 28.286 58.106 20.296 33.752 38.742 58.262 48.601 49.737 5.475 635.96 583.57
mPLUG-Owl2.1 60.295 25.611 33.332 53.647 63.117 47.025 85.635 35.675 73.477 33.926 60.665 55.115 75.605 58.205 60.485 22.193 843.93 784.99
AdapterV2 54.447 25.727 27.113 73.547 70.134 39.1410 86.438 38.348 82.028 33.539 37.329 48.389 76.294 50.407 57.205 7.193 807.18 754.35
VPGTrans 47.003 19.113 26.333 64.323 69.543 46.179 76.224 30.474 76.764 38.276 42.626 44.816 74.149 46.003 56.044 2.509 760.30 714.86
MultiGPT 36.295 23.933 21.007 52.917 62.031 30.262 68.357 25.582 72.767 10.175 58.535 46.355 67.587 49.802 59.824 2.811 687.63 643.24
LaVIN-7B 75.113 25.673 22.892 39.322 40.752 26.847 71.415 25.735 69.977 29.461 37.221 33.061 60.087 50.407 60.487 12.345 680.73 632.17
LaVIN-13B 77.545 27.672 25.227 53.644 48.794 32.774 79.977 27.631 73.547 27.207 35.396 40.686 62.766 49.601 57.587 11.561 731.54 678.65
Lynx 38.287 22.725 19.225 64.327 67.719 42.796 74.777 22.372 73.727 10.282 46.006 47.0010 73.522 51.607 43.967 9.229 707.48 665.54
Fuyu-8B 45.715 27.0610 26.145 48.547 55.467 46.349 83.493 17.103 78.373 40.529 58.822 50.815 74.533 53.001 61.442 1.727 769.05 715.85
LaVIT 42.095 27.223 26.563 61.652 68.742 41.782 73.098 28.952 64.102 15.502 36.841 43.362 65.074 51.406 56.003 0.623 702.97 649.19
Qwen-VL-Chat 63.513 27.836 35.335 62.386 69.066 49.291 85.937 33.926 80.417 44.121 65.487 59.724 73.747 54.401 61.107 20.473 886.69 823.53
BLIP-2 74.171 28.005 35.2210 66.264 68.229 51.063 88.999 43.422 86.886 45.796 70.785 64.423 77.599 58.007 72.022 34.691 965.51 902.29
InstructBLIP 73.332 28.677 33.446 71.606 70.376 52.368 88.689 43.282 85.989 45.689 63.075 62.723 80.5310 58.209 73.107 36.722 967.73 905.62

### v2

@@ 5.1 Best Performance
ChatGPT 69.417 - - 56.553 53.187 48.173 89.605 44.4410 84.975 40.771 65.484 59.9710 72.703 60.848 69.027 38.282 - 783.97
LLaMA1-7B 36.195 - - 67.231 60.721 38.261 82.012 34.262 75.621 15.501 58.533 46.433 65.677 50.404 58.994 2.663 - 656.28
LLaMA1-13B 43.336 - - 66.992 68.828 44.686 72.105 34.112 79.552 28.742 52.073 47.247 65.492 49.205 57.534 19.225 - 685.74
LLaMA2-7B 43.086 - - 66.991 69.222 40.284 67.681 26.381 77.301 16.781 58.533 46.602 67.192 52.001 56.331 3.597 - 648.87
LLaMA2-13B 55.787 - - 66.021 68.697 45.786 81.862 31.496 81.662 24.336 60.375 48.545 69.102 55.002 60.231 20.005 - 713.07
LLaMA3.1-8B-Instruct 60.343 - - 62.1410 61.946 46.081 87.317 34.945 75.127 31.301 50.637 50.811 70.156 58.407 56.043 38.126 - 722.98
Mixtral-AWQ 66.633 - - 54.379 55.5910 43.941 87.925 41.237 79.305 38.277 55.451 60.212 70.561 58.005 64.387 37.039 - 746.25
Gemma 52.955 - - 67.728 61.619 43.102 81.657 36.117 77.052 25.067 54.295 52.435 70.945 57.801 60.072 30.312 - 718.14
Flan-T5-XXL 67.439 - - 64.814 66.014 49.563 89.6010 42.866 86.526 46.296 72.135 63.703 74.398 57.407 71.409 31.419 - 816.08
GPT-4V 79.8010 31.113 57.789 76.195 71.054 50.583 90.913 61.193 87.108 49.441 53.858 60.162 78.953 78.008 76.769 54.298 1057.16 888.47
GPT-4o 87.683 37.789 52.8710 71.439 71.715 53.965 89.399 53.733 86.388 45.254 51.924 57.724 81.382 80.001 70.122 54.298 1045.61 867.28
Claude3-V 76.355 36.1110 53.337 80.954 69.085 46.1510 78.791 32.841 79.937 30.731 38.465 54.475 76.321 66.005 71.372 57.1410 948.02 782.23
GLM-4V-plus 95.073 35.566 57.1410 78.579 70.496 49.425 89.395 51.475 87.817 48.881 56.737 64.237 81.588 66.008 78.427 38.579 1049.33 861.56
Gemini-V 80.355 26.727 47.117 72.733 70.183 51.653 88.348 48.688 87.147 38.867 54.518 59.328 81.023 68.324 56.8310 41.414 973.17 818.99
OpenFlamingo 39.275 23.675 24.447 55.587 61.157 29.479 79.977 24.852 77.307 12.122 57.285 46.195 66.917 49.402 52.681 3.126 703.40 616.02
Fromage 34.517 21.947 20.337 29.856 28.192 22.761 57.197 19.152 47.412 11.042 19.966 27.316 35.106 37.602 40.687 0.161 453.18 376.40
LLaVA-v0-7B 41.105 19.945 24.854 69.423 65.426 30.448 74.692 30.039 74.657 18.129 35.103 44.573 73.483 39.805 43.2110 4.068 688.88 602.99
LLaVA-v0-13B 47.747 12.612 15.187 73.067 69.612 38.514 80.187 30.906 76.587 28.373 37.995 48.465 77.699 40.405 44.295 3.915 725.48 649.95
LLaVA-v1.6-7B 65.895 28.114 34.767 59.952 67.232 45.121 85.639 30.416 81.621 23.977 59.319 52.841 75.189 57.802 59.615 17.191 844.62 715.86
LLaVA-v1.6-13B 69.665 26.899 39.533 64.563 60.433 53.223 86.393 35.969 78.263 40.897 58.737 56.0810 77.9510 61.204 62.313 22.813 894.87 758.79
MiniGPT4 58.705 27.335 30.565 71.125 70.782 50.294 83.994 35.422 83.382 38.465 47.167 49.439 76.012 54.2010 57.493 2.816 837.13 720.54
mPLUG-Owl 37.937 26.173 26.226 51.946 50.366 33.371 68.751 28.286 58.106 20.296 33.752 38.742 58.262 48.601 49.737 5.475 635.96 545.64
mPLUG-Owl2.1 60.295 25.611 33.332 53.647 63.117 47.025 85.635 35.675 73.477 33.926 60.665 55.115 75.605 58.205 60.485 22.193 843.93 724.70
AdapterV2 54.447 25.727 27.113 73.547 70.134 39.1410 86.438 38.348 82.028 33.539 37.329 48.389 76.294 50.407 57.205 7.193 807.18 699.91
VPGTrans 47.003 19.113 26.333 64.323 69.543 46.179 76.224 30.474 76.764 38.276 42.626 44.816 74.149 46.003 56.044 2.509 760.30 667.86
MultiGPT 36.295 23.933 21.007 52.917 62.031 30.262 68.357 25.582 72.767 10.175 58.535 46.355 67.587 49.802 59.824 2.811 687.63 606.95
LaVIN-7B 75.113 25.673 22.892 39.322 40.752 26.847 71.415 25.735 69.977 29.461 37.221 33.061 60.087 50.407 60.487 12.345 680.73 557.06
LaVIN-13B 77.545 27.672 25.227 53.644 48.794 32.774 79.977 27.631 73.547 27.207 35.396 40.686 62.766 49.601 57.587 11.561 731.54 601.11
Lynx 38.287 22.725 19.225 64.327 67.719 42.796 74.777 22.372 73.727 10.282 46.006 47.0010 73.522 51.607 43.967 9.229 707.48 627.26
Fuyu-8B 45.715 27.0610 26.145 48.547 55.467 46.349 83.493 17.103 78.373 40.529 58.822 50.815 74.533 53.001 61.442 1.727 769.05 670.14
LaVIT 42.095 27.223 26.563 61.652 68.742 41.782 73.098 28.952 64.102 15.502 36.841 43.362 65.074 51.406 56.003 0.623 702.97 607.10
LLaMA-3.2-11B-Vision 86.125 26.392 46.115 63.115 63.316 52.085 85.325 38.161 78.807 34.3910 59.507 56.971 77.1810 58.602 57.045 34.6910 917.77 759.15
MiniCPM-V2.6 93.115 28.174 42.273 70.394 70.1810 52.801 89.1410 44.7410 83.636 43.901 44.657 52.847 79.7410 64.4010 67.167 42.665 969.78 806.23
BLIP-2 74.171 28.005 35.2210 66.264 68.229 51.063 88.999 43.422 86.886 45.796 70.785 64.423 77.599 58.007 72.022 34.691 965.51 828.12
InstructBLIP 73.332 28.677 33.446 71.606 70.376 52.368 88.689 43.282 85.989 45.689 63.075 62.723 80.5310 58.209 73.107 36.722 967.73 832.29
GLM-4V-9B 97.035 29.333 43.675 70.875 70.3110 52.563 87.612 40.353 81.377 45.432 52.364 53.654 81.346 63.204 60.527 34.535 964.13 794.10
InternVL2.5-8B 95.145 27.065 47.445 67.486 67.046 49.683 86.095 41.527 80.485 42.362 61.049 59.169 79.311 63.608 67.045 47.445 981.88 812.24
DeepSeek-VL2-tiny 76.655 25.835 35.335 72.827 70.707 45.401 83.643 33.925 83.203 27.629 55.166 53.406 78.346 53.404 65.422 17.037 877.86 740.05
DeepSeek-VL2-small 71.665 26.324 35.223 61.172 60.762 40.5410 80.284 24.852 74.044 39.974 57.867 48.597 76.252 57.604 60.6110 8.915 824.63 691.43
DeepSeek-VL2 70.805 26.446 40.225 68.4510 67.6910 48.015 86.095 36.843 84.022 40.756 65.965 56.893 72.0010 57.203 62.317 29.533 913.2 775.74
Qwen-VL-Chat 63.513 27.836 35.335 62.386 69.066 49.291 85.937 33.926 80.417 44.121 65.487 59.724 73.747 54.401 61.107 20.473 886.69 760.02
Qwen2-VL-7B-Instruct 82.765 30.064 46.115 55.831 61.741 50.927 87.466 41.645 84.271 45.296 65.195 60.375 78.895 66.408 66.217 35.624 958.76 799.83
Qwen2-VL-72B-Instruct 71.945 29.945 44.895 55.106 58.146 50.2610 90.529 49.565 83.559 47.634 68.185 62.481 74.436 61.8010 76.172 52.196 976.78 830.01
Qwen2.5-VL-3B-Instruct 78.785 30.833 42.897 69.903 68.353 51.985 87.923 40.946 78.553 46.326 64.327 59.977 80.189 60.001 62.766 28.283 951.97 799.47
Qwen2.5-VL-7B-Instruct 87.365 31.175 44.913 68.455 66.915 53.773 89.147 42.117 85.565 45.968 64.427 61.266 81.389 64.0010 64.306 44.225 994.92 831.48
We present the best performance ($A^{\tilde{i}}$), calculated using Eq. 2, achieved by different models across various multimodal datasets. This metric serves as a set of benchmarks for evaluating the performance of different models on each dataset. Based on these results, we make the following observations.
5.1.1 Performance of Various Models on Full Datasets
We first report the best performance of various models across full datasets, as detailed in Table 4. Many later models were trained on the ScienceQA dataset, while the MMMU and PuzzleVQA datasets require models to utilize visual context to answer questions. For this reason, we also calculate the total result for each model excluding the VQAMC task (Total).
(1) Comparison between open-source models and closed-source models: Closed-source models such as ChatGPT, GPT-4V, and GPT-4o show excellent performance on various evaluation metrics. Among the open-source alternatives, the Flan-T5 series (including Flan-T5, BLIP-2, and Instruct-BLIP) and Qwen2.5-VL models (such as Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct) show excellent capabilities in multimodal reasoning tasks, but there is still a gap compared to closed-source models with excellent performance. There are two possible reasons for this performance gap. The first reason is training data uncertainty. We cannot be sure whether the closed-source models are trained on our evaluation dataset, which can explain their excellent performance. Another reason is the model size difference. The architecture and parameters of closed-source models may be much larger compared to open-source alternatives. However, we cannot verify this due to limited access to their architecture specifications and training details. (2) Flan-T5 backbone vs. LLaMA backbone in the open-source models: Among open-source models, except for the VQAMA task, InstructBLIP stands out by achieving top performance on three datasets and securing the first position in the ‘Total’ column in the Table 4. Notably, InstructBLIP leverages Flan-T5-XXL as its backbone model. The superior performance of Flan-T5-XXL can be attributed to its Encoder-Decoder architecture, which contrasts with the Decoder-only architecture of the LLaMA series models. The encoder module in Flan-T5-XXL offers significant advantages in multimodal representation learning, making it particularly well-suited for addressing multimodal reasoning tasks with complex multimodal contexts. (3) LLMs vs. MLLMs: For the cumulative best performance of models across all datasets—excluding the VQAMA task, as indicated by the “Total” column—we observe that LLMs exhibit superior performance compared to certain MLLMs, such as Openflamingo, Fromage, LaVIT, and others. The phenomenon may be due to theses MLLMs tend to specialize in traditional visual language (VL) capabilities, including language generation, recognition, knowledge reasoning, spatial reasoning, and OCR, linked to the image modality. Meanwhile, these models might not sufficiently prioritize capabilities like multimodal alignment, multimodal fusion, and etc. Such capabilities are critical for effectively handling multimodal reasoning tasks that involve both vision and text contexts. (4) Different scale checkpoints: In the same model series, larger versions, evidenced by increased parameters such as in LLaMA-7B vs. LLaMA-13B, LLaVA-v0-7B vs. LLaVA-v0-13B, and LaVIN-7B vs. LaVIN-13B, generally outperform their smaller counterparts across most datasets. It suggests that models with more parameters and extensive training are likely to achieve superior performance. (5) Latest models: The updated model iterations, such as the LLaMA, Qwen-VL, mPLUG-OWL, and LLaVA series, demonstrate significant improvements over their predecessors, highlighting how advancements in novel architectures and extended training contribute substantially to enhanced model efficacy. Additionally, the latest models, such as Qwen2.5-VL and MiciCPM-V2.6, showcase exceptional performance. Remarkably, newer models with fewer parameters (e.g., Qwen2.5-VL-7B-Instruct) can match the performance of earlier models with significantly more parameters (e.g., Qwen2-VL-72B-Instruct), underscoring the higher information density of the latest models. (6) Various instructions: On one hand, performance peaks vary within the same model when applied to different datasets, contingent on the instructions used. On the other hand, across various models, the optimal performance on a specific dataset often requires model-specific instructions. Despite its importance, the area of selecting the most suitable instructions for diverse datasets and models is under-researched and represents a promising avenue for future exploration. (7) More challenge tasks: Closed-source models, such as Gemini-V, significantly outperform open-source models on more challenging datasets, including MOSI-7 and MOSEI-7, which involve multiple classification categories, MNRE with a long-tail effect, and complex reasoning datasets like PuzzleVQA and MMMU. These results highlight considerable room for improvement in open-source models when addressing such tasks. Closed-source models demonstrate superior performance on sophisticated tasks like PuzzleVQA and multimodal relation extraction for several technical reasons. 1) Advanced architectural design: These models likely employ more sophisticated attention mechanisms and deeper neural architectures that better handle complex reasoning chains and long-range dependencies. 2) Higher-quality training data: Their training likely involves carefully curated datasets specifically targeting complex reasoning patterns and edge cases. 3) More effective pre-training strategies: They possibly utilize advanced techniques like chain-of-thought prompting and multi-task learning during pre-training to develop stronger reasoning capabilities. 4) Broader knowledge integration: The larger scale of training data and parameters enables better integration of world knowledge, which is crucial for handling long-tail phenomena and multi-step reasoning. 5) Better optimization techniques: They may benefit from more sophisticated training objectives and optimization strategies that specifically enhance performance on complex reasoning tasks. Nevertheless, the latest open-source models have shown remarkable progress in addressing these complex challenges. Notably, the Qwen2.5-VL-7B-Instruct model exhibits competitive performance on sophisticated tasks like PuzzleVQA and MNRE. This progress suggests that open-source models are gradually narrowing the performance gap with closed-source counterparts through improved architectural designs, more effective training strategies, and enhanced multimodal integration capabilities. (8) The inference speed: Notably, models using LLaMA as their backbone exhibit slower inference speeds compared to those based on FLAN-T5 and Qwen architectures. This performance difference stems primarily from LLaMA’s higher parameter density and the lack of specialized inference optimizations that are present in FLAN-T5 and Qwen-based models. (9) Data leakage: Recent multimodal language models have demonstrated remarkable performance on ScienceQA, with MiniCPM, InternVL2.5, and GLM-4V-9B standing out as particularly noteworthy. This superior performance can be attributed to the explicit training of these models on the ScienceQA dataset. In contrast, other models, such as the Qwen-VL series, LLaMA-3.2-Vision-Instruct, and DeepSeek-VL series, lack evidence of similar training, as indicated in their related papers. This observation raises broader concerns about potential data leakage issues, which may not be limited to the ScienceQA dataset. (10) With/Without text context on ScienceQA: ScienceQA primarily relies on the vision context to answer questions. To investigate this further, we conduct comparative experiments excluding textual context, as detailed in the “Text context” section of Figure 3. The results of these experiments are presented in Figure 6. Our observations indicate that while most MLLMs exhibit improved performance on the ScienceQA dataset when textual context is included, the degree of improvement is relatively limited. This finding suggests that the presence or absence of textual context has only a minimal impact on the overall performance of the ScienceQA dataset.
5.1.2 Full datasets vs. Subsets
As discussed in 4.2.2, given the high cost and limited availability of certain closed models, we draw inspiration from Consistently Distributed Sampling (CDS) (Yang et al., 2023c) to tailor our evaluation. Following this, we exclusively evaluate GPT-4V, GPT-4o, Claude3-V, and GLM-4V-plus on these subsets across all tasks, selecting corresponding experimental results from other models for comparison. The comparative analysis, highlighting the best performance metrics between the full dataset and its subset, is illustrated in Figure 7. The MMMU and PuzzleVQA datasets require models to utilize visual context to answer questions, so for these two datasets we only test the performance on MLLMs. Therefore, we do not test MMMU and PuzzleVQA datasets on 9 LLMs. Therefore, we use the “LLMs-Total” metric to calculate the sum of the best accuracies across 14 datasets (excluding PuzzleVQA and MMMU) to measure each LLM’s overall performance. Similarly, the “MLLMs-Total” metric sums up the best performances across all 16 datasets, including PuzzleVQA and MMMU to evaluate MLLM performance.
The subplots “LLMs-Total” and “MLLMs-Total” reveal minimal performance differences between models on full datasets and their subsets. This indicates that our evaluation of models like GPT-4V, GPT-4o, and other closed-models on subsets provides an accurate reflection of their overall performance across tasks. Furthermore, GPT-4o consistently achieves superior performance on most tasks. It excels particularly in challenging datasets with extensive classification categories, such as MOSI-7 and MOSEI-7, datasets with long-tail effects like MNRE, and complex reasoning tasks such as PuzzleVQA and MMMU, where it significantly outperforms other models. Notably, some open-source models, including InstructBLIP, BLIP-2, Qwen2.5-VL, and MiniCPM-V-2.6, also deliver competitive results across many subsets. An additional observation is that larger datasets exhibit less variation in experimental outcomes. This stability underscores the value of using subsets from large datasets, as they tend to provide more representative and reliable results.